OpenStudy (anonymous):

What is the solution for x in the proportion 2/6=x/21

OpenStudy (texaschic101):

do you know how to solve proportions ?

OpenStudy (texaschic101):

2/6 = x/21 now we have to cross multiply (6)(x) = (21)(2) can you finish by solving for x ?

OpenStudy (mokeira):

\[\frac{ 2 }{ 6 }=\frac{ x }{ 21 }\] Step 1: crossmultiply \[6x=2 \times 21\] \[6x=42\] Divide both sides by 6 to get the value of x \[x=7\]

OpenStudy (texaschic101):

or.. you can reduce 2/6 to 1/3 leaving us with : 1/3 = x/21 Proportions are nothing but equivalent fractions. so what is the equivalent fraction of 1/3 with a denominator of 21 ?